OpenStudy (anonymous):

Sam is observing the velocity of a car at different times. After two hours, the velocity of the car is 54 km/h. After four hours, the velocity of the car is 58 km/h. Part A: Write an equation in two variables in the standard form that can be used to describe the velocity of the car at different times. Show your work and define the variables used. (5 points) Part B: How can you graph the equation obtained in Part A for the first six hours? (5 points)

OpenStudy (anonymous):

I have v=50+2t for part A

OpenStudy (pawanyadav):

No That's not correct

OpenStudy (pawanyadav):

Find a from above equation.

OpenStudy (pawanyadav):

Sorry I wrongly typed that It should be 58=54 +a(4) As time is 4 hours

OpenStudy (pawanyadav):

If you got that I start the next part

OpenStudy (pawanyadav):

From there a=1 So equation should be v=54+1(t) Where v is velocity And t is time.